What annual rate of return is earned on a $5,000 investment when it grows to $8,500 in four years? (Do not round intermediate calculations. Round your answer to 2 decimal places.)
Answer #1

We use the formula:
A=P(1+r/100)^n
where
A=future value
P=present value
r=rate of interest
n=time period.

8500=5000*(1+r/100)^4

(8500/5000)^(1/4)=(1+r/100)

(1+r/100)=1.1419

r=1.1419-1

=14.19%(Approx).